# Robust Implicit Regularization via Weight Normalization

### v3

@@ 1 Introduction
${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u};$ (1)
(S)GD is then implemented separately with respect to the vector ${\bf u}$ and magnitude scalar $r$. While overwhelming empirical evidence shows that weight normalization induces faster convergence of (S)GD towards solutions with better generalization performance, rigorous theoretical proofs of these effects have remained challenging due to the nonlinearity introduced in (1).
@@ 1.1 Our contribution and related work
$\Big{\|}{\bf A}\big{(}{\bf x}^{(1)}\odot\cdots\odot{\bf x}^{(L)}\big{)}-{\bf b}\Big{\|}_{2}^{2}$ (2)
where $\odot$ is the entry-wise product. It is shown in [8] that GF on (2) under identical initialization is equivalent to GF on
$\Big{\|}{\bf A}{\bf x}^{\odot L}-{\bf b}\Big{\|}_{2}^{2}.$ (3)
On the other hand, [39] showed that applying WN (1) to overparameterized linear regression induces implicit regularization towards minimal $\ell_{2}$ norm, that is, GF on the loss
$\Big{\|}{\bf A}\Big{(}\frac{r}{\|{\bf u}\|_{2}}{\bf u}\Big{)}-{\bf b}\Big{\|}_{2}^{2}$ (4)
@@
$\lim_{t\to\infty}r(t){\bf u}(t)\approx\operatorname*{arg\,min}_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{2}.$ (5)
@@
$\lim_{t\to\infty}(r(t){\bf u}(t))^{\odot L}\approx\operatorname*{arg\,min}_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}.$ (7)
@@
(a) Weight normalization significantly decreases the error compared to gradient descent, especially from larger-scale initialization of the weights.
(b) WN enables algorithms to have small error even without small initialization.
Figure 1: WN converges to minimal $\ell_{1}$-norm solutions from a wider range of initialization, and hence is more robust in the sense that it is not sensitive to the choice of initialization. This suggests that GF with WN could be used as an efficient alternative for $\ell_{1}$-minimization. Each data point in Figure 0(a) is an average over ten random initializations at fixed scale. In Figure 0(a), the plotted errors are after a fixed common number of iterations of GF with and without WN, so larger error indicates a combination of larger final error and slower convergence.
Our work also builds on a long line of work on the factorized linear multilayer model [21, 24, 37, 38, 42, 43]. To precisely quantify the error coming from the approximation $\approx$” in (5), we seek inspiration from [8], where the authors showed that for $L\geq 2$, GF on the loss (3) converges to the limit whose $\ell_{1}$-norm is almost minimized in the following sense. Let $\tilde{{\bf x}}={\bf x}^{\odot L}$. Then the limit $\tilde{{\bf x}}_{\infty}:=\lim_{t\to\infty}\tilde{{\bf x}}(t)$ satisfies ${\bf A}\tilde{{\bf x}}_{\infty}={\bf b}$ and
$\|\tilde{{\bf x}}_{\infty}\|_{1}-\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}\leq\varepsilon(\|\tilde{{\bf x}}_{0}\|_{1})\cdot\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}$ (8)
where $\tilde{{\bf x}}_{0}=\tilde{{\bf x}}(0)$ and $\varepsilon:\mathbb{R}\to\mathbb{R}$ is an increasing function (which we will later used and specified) with $\varepsilon(0)=0$. In general, we need $\|\tilde{{\bf x}}_{0}\|_{1}\ll 1$ to have $\varepsilon(\|\tilde{{\bf x}}_{0}\|_{1})\ll 1$. This often implies that if the initialization is on the unit sphere, the error $\varepsilon$ is simply too large for (8) to be useful at all.
Fortunately, this problem disappears when we incorporate normalization as in (6). In Theorem 2.2, we show that the error bound can be improved to
$\|\tilde{{\bf x}}_{\infty}\|_{1}-\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}\leq\varepsilon(\rho^{-L}\|\tilde{{\bf x}}_{0}\|_{1})\cdot\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}$ (9)
for some $\rho\geq 1$ that increases as ${\bf x}_{0}$ decreases (the explicit dependence will be later stated in (18)). As a consequence, the right hand side remains small even for moderately small initialization such as the normalized Gaussian vector, which often leads to smaller generalization error [11, 23].
@@ 1.2 Formulation
$\mathcal{L}({\bf x})$ $=\frac{1}{2L}\|{\bf A}{\bf x}^{\odot L}-{\bf b}\|_{2}^{2}.$ (10)
@@
$\partial_{t}{\bf x}$ $=-\nabla\mathcal{L}({\bf x}),\quad{\bf x}(0)={\bf x}_{0}.$ (11)
@@
~$\widetilde{\mathcal{L}}(r,{\bf u})$ $=\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right)$ (12)
@@
$\partial_{t}r$ $=-\eta_{r}\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u}),\quad r(0)=r_{0}$ (13)
$\partial_{t}{\bf u}$ $=-\eta_{\bf u}\nabla_{{\bf u}}\widetilde{\mathcal{L}}(r,{\bf u}),\quad{\bf u}(0)={\bf u}_{0}$ (14)
@@ 2 Main Results
The core idea is to magnify the implicit regularization via an appropriate learning rate ratio to obtain small error without small initialization $r_{0}$. For instance, we will see that in the setting of Theorem 2.2, the error decreases exponentially with respect to initialization according to (20), which can be compared to the setting of [8] where the error only decreases polynomially according to (8).
@@
Let $L\geq 2$, ${\bf A}\in\mathbb{R}^{M\times N}$ and ${\bf b}\in\mathbb{R}^{M}$ and assume that $S_{+}$ is non-empty. Suppose ${\bf x}$ follows the dynamics (11) with ${\bf x}_{0}>0$. Let $\tilde{{\bf x}}={\bf x}^{\odot L}$.Then the limit $\tilde{{\bf x}}_{\infty}:=\lim_{t\to\infty}\tilde{{\bf x}}(t)$ exists and $\tilde{{\bf x}}_{\infty}\in S_{+}$. Moreover, consider $\gamma=1-\frac{2}{L}\in[0,1]$ and
@@
$\|\tilde{{\bf x}}_{\infty}\|_{{\bf w},1}-{Q}\leq\varepsilon{Q},$ (15)
@@
$c_{L}:=\begin{cases}1&\text{if }L=2,\\ \left(\frac{L}{2}\right)^{\frac{L}{L-2}}&\text{if }L>2,\end{cases}$ (16)
@@
$\varepsilon(\beta_{1},\beta_{\min}):=\begin{cases}\frac{\log(\beta_{1}/\beta_{\min})}{\log({Q}/\beta_{1})}&\text{if }L=2,\\[6.0pt] \frac{L(\beta_{1}^{\gamma}-\beta_{\min}^{\gamma})}{2{Q}^{\gamma}-L\beta_{1}^{\gamma}}&\text{if }L>2.\end{cases}$ (17)
@@
Let $L\in\mathbb{N}$, $L\geq 2$, ${\bf A}\in\mathbb{R}^{M\times N}$, ${\bf b}\in\mathbb{R}^{M}$, $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$ for some constant ${\tilde{\eta}}>0$. Suppose $(r,{\bf u})$ follow the dynamics in (13) and (14) with $r_{0},{\bf u}_{0}>0$ satisfying $\|{\bf u}_{0}\|_{2}=1$ and $r_{0}\leq\min({\tilde{\eta}}^{1/2},\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L})$. Denote ${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Suppose there exists ${\bf v}>0$ such that ${\bf A}{\bf v}=0$, $S_{+}$ is non-empty, and the limit $\tilde{{\bf x}}_{\infty}:=\lim_{t\to\infty}{\bf x}(t)^{\odot L}$ exists. Let
$\rho:=\frac{r_{0}}{\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}}\exp\left(\frac{\|{\bf A}^{\dagger}{\bf b}\|_{2}^{2/L}-r_{0}^{2}}{2{\tilde{\eta}}}\right),$ (18)
@@
    The loss defined in (10) decreases exponentially in time, i.e., for all $t\geq 0$
    $\mathcal{L}({\bf x}(t))\leq\mathcal{L}({\bf x}_{0})e^{-ct}$ (19)
@@
    It holds that $\rho\geq 1$, and the limit $\tilde{{\bf x}}_{\infty}$ satisfies (15) with error
    $\varepsilon(\rho^{-L}\beta_{1},\rho^{-L}\beta_{\min})$ (20)
    as defined in (17).
@@
  • How to choose ${\tilde{\eta}}$ and $r_{0}$.
    Since the goal is to make $\rho$ large, we want to set ${\tilde{\eta}}$ small. However, due to the constraint $r_{0}\leq{\tilde{\eta}}^{1/2}$, small ${\tilde{\eta}}$ will also restrict the range of $r_{0}$ we can choose. Hence there is a trade-off. Fortunately, since ${\tilde{\eta}}$ contributes an exponential factor, mildly small ${\tilde{\eta}}$ like $0.1$ is usually sufficient.
@@
    ±$\mathcal{L}_{\pm}({\bf u},{\bf v})=\|{\bf A}({\bf u}^{\odot L}-{\bf v}^{\odot L})-{\bf b}\|_{2}^{2},\quad{\bf u}_{0},{\bf v}_{0}>0$ (21)
@@
    for any $\zeta$. In particular, since we cannot assume ${\bf u}$ and ${\bf v}$ have disjoint supports, it is unlikely that we can uniquely define $({\bf u},{\bf v})$ based on the invariants in Lemma 3.8. Therefore, our analysis only focuses on the case of positive solution set in this paper. We nevertheless include some simulation for gradient descent on the modified loss function (21) in Section 4.3 along with some discussion.
Theorem 2.2 provides the optimal scaling when $\eta_{r}$ is a constant. Interestingly, an alternative time-dependent step-size choice $(\eta_{r},\eta_{\bf u})=(r^{2},1)$ provides a different dynamic which is close to the gradient flow dynamics without weight normalization (11). In this case, we can prove convergence to the stationary point instead of assuming it for all $L\geq 2$. Below we state the results for this special case and encourage readers to explore more possibilities.
@@
    The limit $\tilde{{\bf x}}_{\infty}$ satisfies (15) with error $\varepsilon(\beta_{1},\beta_{\min})$ as defined in (17). In other words, it satisfies the same error bound as in Theorem 2.2 with $\rho=1$ (no magnification).

### v4

@@ 1 Introduction
${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u};$ (1.1)
(S)GD is then implemented separately with respect to the vector ${\bf u}$ and magnitude scalar $r$. While overwhelming empirical evidence shows that weight normalization induces faster convergence of (S)GD towards solutions with better generalization performance, rigorous theoretical proofs of these effects have remained challenging due to the nonlinearity introduced in (1.1).
@@ 1.1 Our contribution and related work
$\Big{\|}{\bf A}\big{(}{\bf x}^{(1)}\odot\cdots\odot{\bf x}^{(L)}\big{)}-{\bf b}\Big{\|}_{2}^{2}$ (1.2)
where $\odot$ is the entry-wise product. It is shown in [8] that GF on (1.2) under identical initialization is equivalent to GF on
$\Big{\|}{\bf A}{\bf x}^{\odot L}-{\bf b}\Big{\|}_{2}^{2}.$ (1.3)
On the other hand, [39] showed that applying WN (1.1) to overparameterized linear regression induces implicit regularization towards minimal $\ell_{2}$ norm, that is, GF on the loss
$\Big{\|}{\bf A}\Big{(}\frac{r}{\|{\bf u}\|_{2}}{\bf u}\Big{)}-{\bf b}\Big{\|}_{2}^{2}$ (1.4)
@@
$\lim_{t\to\infty}r(t){\bf u}(t)\approx\operatorname*{arg\,min}_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{2}.$ (1.5)
@@
$\lim_{t\to\infty}(r(t){\bf u}(t))^{\odot L}\approx\operatorname*{arg\,min}_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}.$ (1.7)
@@
(a) Gradient descent with weight normalization (WN-GD) yields significantly smaller reconstruction error.
(b) For a fixed error upper bound, WN-GD allows a much wider range of initialization. Here $\delta$ is given in (2.7).
Figure 1: WN converges to minimal $\ell_{1}$-norm solutions from a wider range of initialization, and hence is more robust in the sense that it is not sensitive to the choice of initialization. This suggests that GF with WN could be used as an efficient alternative for $\ell_{1}$-minimization. Each data point in Figure 1(a) is an average over multiple random initializations at fixed scale. The improvement ratio (reconstruction error for GD divided by reconstruction error for WN-GD) can be huge; when the initialization scale $\alpha$ (defined in (4.1)) equals to $0.1$, such ratio is more than $10^{5}$!
Our work also builds on a long line of work on the factorized linear multilayer model [21, 24, 37, 38, 42, 43]. To precisely quantify the error coming from the approximation $\approx$” in (1.5), we seek inspiration from [8], where the authors showed that for $L\geq 2$, GF on the loss (1.3) converges to the limit whose $\ell_{1}$-norm is almost minimized in the following sense. Let $\tilde{{\bf x}}={\bf x}^{\odot L}$. Then the limit $\tilde{{\bf x}}_{\infty}:=\lim_{t\to\infty}\tilde{{\bf x}}(t)$ satisfies ${\bf A}\tilde{{\bf x}}_{\infty}={\bf b}$ and
$\|\tilde{{\bf x}}_{\infty}\|_{1}-\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}\leq\varepsilon(\|\tilde{{\bf x}}_{0}\|_{1})\cdot\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}$ (1.8)
where $\tilde{{\bf x}}_{0}=\tilde{{\bf x}}(0)$ and $\varepsilon:\mathbb{R}\to\mathbb{R}$ is an increasing function (which we will later used and specified) with $\varepsilon(0)=0$. In general, we need $\|\tilde{{\bf x}}_{0}\|_{1}\ll 1$ to have $\varepsilon(\|\tilde{{\bf x}}_{0}\|_{1})\ll 1$. This often implies that if the initialization is on the unit sphere, the error $\varepsilon$ is simply too large for (1.8) to be useful at all.
Fortunately, this problem disappears when we incorporate normalization as in (1.6). In Theorem 2.2, we show that the error bound can be improved to
$\|\tilde{{\bf x}}_{\infty}\|_{1}-\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}\leq\varepsilon(\rho^{-L}\|\tilde{{\bf x}}_{0}\|_{1})\cdot\min_{{\bf A}{\bf z}={\bf b}}\|{\bf z}\|_{1}$ (1.9)
for some $\rho\geq 1$ that increases as ${\bf x}_{0}$ decreases (the explicit dependence will be later stated in (2.4)). As a consequence, the right hand side remains small even for moderately small initialization such as the normalized Gaussian vector, which often leads to smaller generalization error [11, 23].
@@ 1.2 Formulation
$\mathcal{L}({\bf x})$ $=\frac{1}{2L}\|{\bf A}{\bf x}^{\odot L}-{\bf b}\|_{2}^{2}.$ (1.10)
@@
$\partial_{t}{\bf x}$ $=-\nabla\mathcal{L}({\bf x}),\quad{\bf x}(0)={\bf x}_{0}.$ (1.11)
@@
~$\widetilde{\mathcal{L}}(r,{\bf u})$ $=\mathcal{L}\left(\frac{r}{\|{\bf u}\|_{2}}{\bf u}\right)$ (1.12)
@@
$\partial_{t}r$ $=-\eta_{r}\nabla_{r}\widetilde{\mathcal{L}}(r,{\bf u}),\quad r(0)=r_{0}$ (1.13)
$\partial_{t}{\bf u}$ $=-\eta_{\bf u}\nabla_{{\bf u}}\widetilde{\mathcal{L}}(r,{\bf u}),\quad{\bf u}(0)={\bf u}_{0}$ (1.14)
@@ 2 Main Results
The core idea is to magnify the implicit regularization via an appropriate learning rate ratio to obtain small error without small initialization $r_{0}$. For instance, we will see that in the setting of Theorem 2.2, the error decreases exponentially with respect to initialization according to (2.6), which can be compared to the setting of [8] where the error only decreases polynomially according to (1.8).
@@
Let $L\geq 2$, ${\bf A}\in\mathbb{R}^{M\times N}$ and ${\bf b}\in\mathbb{R}^{M}$ and assume that $S_{+}$ is non-empty. Suppose ${\bf x}$ follows the dynamics (1.11) with ${\bf x}_{0}>0$. Let $\tilde{{\bf x}}={\bf x}^{\odot L}$. Then the limit $\tilde{{\bf x}}_{\infty}:=\lim_{t\to\infty}\tilde{{\bf x}}(t)$ exists and $\tilde{{\bf x}}_{\infty}\in S_{+}$. Moreover, let
@@
$\|\tilde{{\bf x}}_{\infty}\|_{{\bf w},1}-{Q}\leq\varepsilon{Q},$ (2.1)
@@
$c_{L}:=\begin{cases}1&\text{if }L=2,\\ \left(\frac{L}{2}\right)^{\frac{L}{L-2}}&\text{if }L>2,\end{cases}$ (2.2)
@@
$\varepsilon(\beta_{1},\beta_{\min}):=\begin{cases}\frac{\log(\beta_{1}/\beta_{\min})}{\log({Q}/\beta_{1})}&\text{if }L=2,\\[6.0pt] \frac{L(\beta_{1}^{1-L/2}-\beta_{\min}^{1-L/2})}{2{Q}^{1-L/2}-L\beta_{1}^{1-L/2}}&\text{if }L>2.\end{cases}$ (2.3)
@@
Let $L\in\mathbb{N}$, $L\geq 2$, ${\bf A}\in\mathbb{R}^{M\times N}$, ${\bf b}\in\mathbb{R}^{M}$, $(\eta_{r},\eta_{\bf u})=({\tilde{\eta}},1)$ for some constant ${\tilde{\eta}}>0$. Suppose $(r,{\bf u})$ follow the dynamics in (1.13) and (1.14) with $r_{0},{\bf u}_{0}>0$ satisfying $\|{\bf u}_{0}\|_{2}=1$ and ${\tilde{\eta}}^{1/2}\leq r_{0}\leq\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}$. Denote ${\bf x}=\frac{r}{\|{\bf u}\|_{2}}{\bf u}$. Suppose there exists ${\bf v}>0$ such that ${\bf A}{\bf v}=0$, $S_{+}$ is non-empty, and the limit $\tilde{{\bf x}}_{\infty}:=\lim_{t\to\infty}{\bf x}(t)^{\odot L}$ exists. Let
$\rho:=\frac{r_{0}}{\|{\bf A}^{\dagger}{\bf b}\|_{2}^{1/L}}\exp\left(\frac{\|{\bf A}^{\dagger}{\bf b}\|_{2}^{2/L}-r_{0}^{2}}{2{\tilde{\eta}}}\right),$ (2.4)
@@
    The loss defined in (1.10) decreases exponentially in time, i.e., for all $t\geq 0$
    $\mathcal{L}({\bf x}(t))\leq\mathcal{L}({\bf x}_{0})e^{-ct}$ (2.5)
@@
    It holds that $\rho\geq 1$, and the limit $\tilde{{\bf x}}_{\infty}$ satisfies (2.1) with error
    $\varepsilon(\rho^{-L}\beta_{1},\rho^{-L}\beta_{\min})$ (2.6)
    as defined in (2.3).
@@
  • Enlarged range of initialization for sparse recovery.
    Based on the error reduction, below we show that there is a wider range of initialization that yields similar (or even smaller error) for weight normalized GF than the regular GF. The ratio is given by $\|{\bf A}^{\dagger}{\bf b}\|_{2}/\delta$, where $\delta\ll 1$ is the small radius for GD to achieve sufficient implicit regularization. From (1.8) and (1.9), we know that the equivalent level sets for errors are given by
    $\begin{cases}&\text{GD: }\{\tilde{{\bf x}}_{0}:\|\tilde{{\bf x}}_{0}\|_{1}\leq\delta\}\\ &\text{WN-GD: }\{(r_{0}{\bf u}_{0})^{\odot L}:\rho^{-L}\|(r_{0}{\bf u}_{0})^{\odot L}\|_{1}\leq\delta\}.\end{cases}$ (2.7)
    Consider $L=2$. Since $\|{\bf u}_{0}\|_{2}=1$, if ${\bf z}=(r_{0}{\bf u}_{0})^{\odot 2}$, then $\|{\bf z}\|_{1}=\|(r_{0}{\bf u}_{0})^{\odot 2}\|_{1}=r_{0}^{2}$. Hence
    $\{(r_{0}{\bf u}_{0})^{\odot 2}:\rho^{-2}\|(r_{0}{\bf u}_{0})^{\odot 2}\|_{1}\leq\delta\}$
    =$\qquad=\left\{(r_{0}{\bf u}_{0})^{\odot 2}:\frac{\|{\bf A}^{\dagger}{\bf b}\|_{2}}{r_{0}^{2}}\exp\left(\frac{r_{0}^{2}-\|{\bf A}^{\dagger}{\bf b}\|_{2}}{{\tilde{\eta}}}\right)\|(r_{0}{\bf u}_{0})^{\odot 2}\|_{1}\leq\delta\right\}$
    =$\qquad=\left\{{\bf z}:\frac{\|{\bf A}^{\dagger}{\bf b}\|_{2}}{\|{\bf z}\|_{1}}\exp\left(\frac{\|{\bf z}\|_{1}-\|{\bf A}^{\dagger}{\bf b}\|_{2}}{{\tilde{\eta}}}\right)\|{\bf z}\|_{1}\leq\delta\right\}$
    =$\qquad=\left\{{\bf z}:\exp\left(\frac{\|{\bf z}\|_{1}-\|{\bf A}^{\dagger}{\bf b}\|_{2}}{{\tilde{\eta}}}\right)\leq\frac{\delta}{\|{\bf A}^{\dagger}{\bf b}\|_{2}}\right\}$
    =$\qquad=\{{\bf z}:\|{\bf z}\|_{1}\leq\|{\bf A}^{\dagger}{\bf b}\|_{2}+{\tilde{\eta}}(\log\delta-\log(\|{\bf A}^{\dagger}{\bf b}\|_{2}))\}$
    Hence the radius of the $\ell_{1}$-ball increases from $\delta$ to $\|{\bf A}^{\dagger}{\bf b}\|_{2}+{\tilde{\eta}}(\log\delta-\log(\|{\bf A}^{\dagger}{\bf b}\|_{2}))$, which leads to a huge improvement when $\delta\ll\|{\bf A}^{\dagger}{\bf b}\|_{2}$. Note that this is often the case because $\delta$ is usually chosen to be small to have small error $\varepsilon$, whereas $\|{\bf A}^{\dagger}{\bf b}\|_{2}$ is not small in general.
@@
    ±$\mathcal{L}_{\pm}({\bf u},{\bf v})=\|{\bf A}({\bf u}^{\odot L}-{\bf v}^{\odot L})-{\bf b}\|_{2}^{2},\quad{\bf u}_{0},{\bf v}_{0}>0$ (2.8)
@@
    for any $\zeta$. In particular, since we cannot assume ${\bf u}$ and ${\bf v}$ have disjoint supports, it is unlikely that we can uniquely define $({\bf u},{\bf v})$ based on the invariants in Lemma 3.8. Therefore, our analysis only focuses on the case of positive solution set in this paper. We nevertheless include some simulation for gradient descent on the modified loss function (2.8) in Section 4.4 along with some discussion.
Theorem 2.2 provides the optimal scaling when $\eta_{r}$ is a constant. Interestingly, an alternative time-dependent step-size choice $(\eta_{r},\eta_{\bf u})=(r^{2},1)$ provides a different dynamic which is close to the gradient flow dynamics without weight normalization (1.11). In this case, we can prove convergence to the stationary point instead of assuming it for all $L\geq 2$. Below we state the results for this special case and encourage readers to explore more possibilities.
@@
    The limit $\tilde{{\bf x}}_{\infty}$ satisfies (2.1) with error $\varepsilon(\beta_{1},\beta_{\min})$ as defined in (2.3). In other words, it satisfies the same error bound as in Theorem 2.2 with $\rho=1$ (no magnification).